# When Attention Collapses: How Degenerate Layers in LLMs Enable Smaller, Stronger Models

### v3

@@ Abstract
Large Language Models (LLMs) leverage the transformer architecture and its self-attention mechanism to achieve remarkable performance. However, we identify a structural inefficiency in standard pre-trained decoder-style LLMs: attention matrices in many deeper layers frequently degenerate, often collapsing to near rank-one, single-column structures. These layers, which we term lazy layers, are redundant and inefficient.
To address this, we introduce $\mathsf{Inheritune}$, a simple yet effective training recipe for developing smaller, more attentive, and high-performing language models. $\mathsf{Inheritune}$ initializes a compact model by inheriting the potent early layers from a larger pre-trained model, then progressively retrains and expands it. Our demonstration of $\mathsf{Inheritune}$’s efficacy on various models and datasets reveals that models trained with this approach, despite having significantly fewer layers, can achieve performance comparable to or even exceeding their larger counterparts. This approach enables model compression by design, producing language models that are both compact and high-performing. Code is available at https://github.com/sanyalsunny111/LLM-Inheritune.
@@ 1 Introduction
Large Language Models (LLMs) are built with decoder-style transformer blocks (Vaswani et al., 2017). These models are typically designed to be large, with a significant portion of their parameters dedicated to their depth, with multiple transformer blocks stacked with eachother building model capacity. Each block or layer in the stack refines the representations learned by the previous blocks, allowing the model to develop a nuanced understanding of the input data. As these models scale in depth and size, their performance tends to improve Kaplan et al. (2020); Hoffmann et al. (2022), benefiting from increased model capacity.
The causal self-attention (hereafter referred to as attention) mechanism is arguably the most crucial component of a transformer block. It allows models to combine tokens as a weighted linear sum of their attention scores, effectively capturing long-range dependencies and contextual relationships within text data. However, as models grow in depth, they often encounter a phenomenon known as attention degeneration caused by collapse in the attention rank ((Noci et al., 2022; Dong et al., 2021; He et al., 2023)). Notably, this phenomenon has not been studied in the context of standard LLMs. A formal discussion on attention degeneration is provided in Section 2.
In this paper, we empirically analyze 24-layer GPT-2 medium and 36-layer GPT-2 large models (decoder-style LLMs) Radford et al. (2019) for attention degeneration and observe that many deeper layers in both models exhibit rank-1 attention matrices. Further investigation reveals that most of these rank-1 matrices are also single-column, i.e. their mass is concentrated to a single column. Our attention matrix analysis is presented in Figure 1. We term these deeper layers, where all attention matrices of a given layer are degenerated, as lazy layers.
@@
Motivated by this new finding we aim to develop performant small base language models (LMs) utilizing weights from inefficient larger base LMs. A base LM is a decoder-style model trained solely for next-token prediction without additional enhancements like instruction tuning or reinforcement learning with human feedback (RLHF). Our proposal is straightforward, we start by initializing our smaller LM (target) using the first few blocks from a large pre-trained LM (reference). We then train the target model for a specified number of steps. After this initial training, we incrementally grow the target model by adding more blocks, continuing the training process until it matches or surpasses the pre-train validation loss (also val loss) of the reference model. During the growth phase, the newly added blocks can be initialized with lazy layers of the reference LM. We refer to this simple yet effective training approach as $\mathsf{Inheritune}$.
(a) GPT-2 Medium (355M) L=24, H=1024
(b) GPT-2 Large (770M) L=36, H=1280
(c) GPT-2 XLarge (1.5B) L=48, H=1600
@@
    Analysis of Attention Degeneration Leading to Lazy Layers. We empirically investigate attention degeneration in standard LLM settings. Our analysis shows that rank-collapsed attention matrices often exhibit single-column structures, revealing a significant structural inefficiency in the attention mechanism of standard LLMs in deeper layers (refer Figure 1).
@@
    Introduction of Inheritune. We propose $\mathsf{Inheritune}$ as an approach to effectively train high-performing, smaller models. This method involves inheriting a few early blocks from a larger pre-trained model and progressively growing and training the smaller model. The initialization is entirely zero-shot. We validate the effectiveness of $\mathsf{Inheritune}$ through comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M), and GPT-2 medium (355M) models, trained on the OpenWebText and the FineWeb_edu dataset.
@@
    Evaluation Against Multiple Baselines. Models derived using $\mathsf{Inheritune}$ consistently outperform various baselines, including much larger models trained from scratch (refer Figure 5 and Figure 6), model initialization and efficient training baselines (refer Table 1), and models trained using two knowledge distillation techniques (refer Figure 7).
@@ 2 Attention Collapse and the Emergence of Lazy Layers in LLMs
A vanilla transformer-based model consists of $L$ transformer blocks (layers). The model operates on an input sequence $X\in\mathbb{R}^{T\times e}$, where $T$ denotes the sequence length (number of tokens), and $e$ represents the embedding dimension or model hidden size. The output of each layer $l$ is denoted as $X^{(l)}\in\mathbb{R}^{T\times e}$.
Each transformer block primarily consists of two sub-layers: a self-attention block and a position-wise feed-forward network (FFN). The self-attention mechanism enables the model to weight the relevance of different tokens in the sequence relative to each other. Specifically, for a single attention head, the attention computation is defined as $\text{Attention}(Q,K,V)=\underbrace{\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)}_{\textbf{Attention matrix: $A(X)$}}V$
@@
where $\tau\in(0,1)$ represents the proportion of variance that must be captured by the top $k$ singular values. A lower value of $k^{*}$ indicates stronger rank collapse. In this work, we set $\tau=0.90$. A lower $k^{*}$ indicates stronger rank collapse.
In Figure 1, we present the layer-wise analysis of rank of GPT-2 models. For this analysis, we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText with 4M tokens, each with a sequence length of $T=100$ tokens across all attention heads within each layer. We then define the average approximate rank for each head and layer as $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$. Subsequently, we aggregate this metric per layer by taking the maximum rank across heads: $\mathrm{MaxRank}^{(l)}=\max_{h}\{\mathrm{Rank}^{(h,l)}\}$. As demonstrated in Figure 1, $\text{MaxRank}^{(l)}$ reveals that many deeper layers exhibit attention matrices that are predominantly near rank-1. We highlight that this rank collapse occurs in GPT-2 medium, Large, and xLarge models—which are standard architectures incorporating skip connections and FFNs—thereby extending the findings of Dong et al. (2021) and Noci et al. (2022) to these widely-used LLMs. The rank analysis is quite robust to the dataset in Figure 10, we performed the same analysis using the same models on held out data from Fineweb_edu dataset (Penedo et al., 2024). Overall, the degeneration of attention matrices in deeper layers provides quantitative evidence for the existence of lazy layers. Specifically, we observe that some deeper layers exhibit a near-complete rank collapse of attention matrices across all heads, suggesting potentially reduced representational capacity and less effective token mixing in these layers.
@@ 2.1 The Functional Ineffectiveness of Lazy Layers
(a) Rank profile (GPT-2 Small 12L).
(b) Performance of 4L variants.
@@
Figure 2: Higher-rank layers transfer better. (Left, 2(a)) Layer-wise $\mathrm{MaxRank}^{(l)}$ of a pre-trained 12L GPT-2 Small. (Right, 2(b)) Validation loss of 4L variants initialized with potent blocks (AvgRank $\approx 8.4-9.5$) vs. a lazy block (AvgRank $\approx 1.2$) or random weights, after 100K steps. Lazy block initialization mirrors random.
@@
In the first set of experiments, we trained a vanilla GPT2-small (125M) model with 12 layers for 100K steps on the OpenWebText dataset. We then performed the rank analysis described earlier, with results presented in Figure 2. Specifically, we aggregated the approximate ranks over groups of contiguous layers using $\text{AvgRank}=\frac{1}{L}\sum_{l=1}^{L}\text{MaxRank}^{(l)}$, where $L$ is the number of layers in each group. Subsequently, we trained three GPT2-small variants [^1] A variant shares the same configurations as the parent model but has fewer layers. for 100K steps, each initialized with a different contiguous block of four layers from the trained vanilla GPT2-small model: (a) layers 1–4, with $\text{AvgRank}=8.40$; (b) layers 5–8, with $\text{AvgRank}=9.48$; and (c) layers 9–12, with $\text{AvgRank}=1.22$. The last model is initialized with lazy layers. For comparison, we also trained another GPT2-small variant with random initialization for 100K steps. All models were trained on the OpenWebText dataset. As shown in Figure 2, the model initialized with layers from the vanilla GPT2-small model having higher AvgRank demonstrated the best performance (i.e., lowest final validation loss). Additionally, we observed that the model initialized with lazy layers performed very similarly to the model with random initialization suggesting that lazy layers contain minimal transferable knowledge. The key results are also summarized in Table 4.
(a) GPT2-medium (12-layer) variants.
(b) GPT2-large (16-layer) variants.
@@
Figure 3: When initializing 12-layer and 16-layer variants of GPT2-medium and GPT2-large with deeper (lazy) layers showing degenerated attention, performance is comparable to random initialization. In contrast, early-layer initialization leads to significantly better convergence and generalization.
@@
For the second set of experiments we utilized larger models namely GPT-2 medium and GPT-2 large both similarly trained for 100K steps using OpenWebText dataset. Here we initialized a 12-layer GPT-2 medium variant and an 18-layer variant of GPT-2 large using lazy layers extracted from pre-trained 24-layer GPT-2 medium and 36-layer GPT-2 large models. We then trained these GPT-2 variants on the same dataset for 10K steps. For comparison, we conducted two baseline experiments where the GPT-2 variants were initialized either with the first half of transformer layers (potent layers with high AvgRank) and with random initialization. As shown in Figure 3, models initialized with lazy layers demonstrate poor transferability, performing similarly to models with random initialization. This provide additional evidence that lazy layers with fully degenerated attention, fails to learn meaningful representations.
@@ 3 Inheritune: Our Proposed Training Recipe
  1. 1.
  2. 2.
  3. 3.
    Grow: If needed, increase $\mathcal{M}_{\text{tgt}}$’s size by adding next few contiguous layers and repeat steps 1-2 until desired performance is achieved.
@@
6:while $\mathcal{M}_{\text{tgt}}$ performance < $\mathcal{M}_{\text{ref}}$ performance on $\mathcal{D}_{\text{val}}$ do
@@
Figure 4: Overview of the $\mathsf{Inheritune}$ training recipe using a 24-Layer GPT-2 medium model example. A smaller target model is initialized using early layers from a larger, pre-trained reference model. The target model goes multiple rounds of training while inheriting more early layers until it matches the reference model. The intensity of the red color in layers correlates with $\mathrm{MaxRank}^{(l)}$.
@@ 4 Experiments
We evaluate $\mathsf{Inheritune}$ through a comprehensive set of experiments using several GPT-2 models: 48-layer GPT-2 xlarge (1.5B), 36-layer GPT-2 large (770M), 32-layer GPT-2 large (668M) and 24-layer GPT-2 medium (355M) models Radford et al. (2019). Table 11 provides detailed specifications of all model configurations used.
We use two training datasets in our experiments: OpenWebText Gokaslan & Cohen (2019) and Fineweb_edu Penedo et al. (2024). Our experimental setup closely follows prior works Liu et al. (2023); Sanyal et al. (2024). For models trained on OpenWebText, we report validation loss (log perplexity) on a held-out set. For those trained on Fineweb-Edu, we report training loss (also log perplexity). In addition, we conduct zero-shot downstream evaluations using the lm-evaluation-harness Gao et al. (2024) across five standard tasks: ARC-easy (ARCE; Clark et al. (2018)), LAMBADA Paperno et al. (2016), SciQ Welbl et al. (2017), HellaSwag Zellers et al. (2019), and PIQA Bisk et al. (2020). We also perform a thorough ablation study of our initialization strategy, focusing on 16-layer GPT-2 medium variant.
We provide experimental details our $\mathsf{Inheritune}$ training recipe using a GPT-2 medium model as an example; similar procedure was applied to train other models. A visualization of the training recipe is presented in Figure 4. Our recipe for applying $\mathsf{Inheritune}$ involves the following steps.
@@
    Reference Model: We train the full 24-layer GPT-2 medium model on $\mathcal{D}_{\text{train}}$ for 100K steps and evaluate its validation loss ( log-perplexity) on $\mathcal{D}_{\text{val}}$. This establishes our benchmark validation loss.
@@
    Iterative Refinement: If the smaller model’s performance is inferior to the reference model, then we incrementally increase its size by two layers and repeat steps 2-3 until we achieve parity with the reference model’s validation loss.
@@
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
Figure 5: Models derived using Inheritune converge faster and match the final validation loss of the full-sized model, despite having much fewer layers. Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge variant, 18-layer GPT-2 Large variant, 16-layer GPT-2 Medium variant) against their full-sized counterparts and same sized variants trained from scratch. All models are trained for 100K steps using OpenWebText data with data repetition.
(a) GPT2-Medium
(b) GPT2-Large
Figure 6: Models derived using $\mathsf{Inheritune}$ converge faster and match the final validation loss and average downstream performance (Perf.) across 5 datasets of the full-sized model, while outperforming same-sized models. (Plots) Validation loss curves for our 16-layer variants versus full-sized models. (Table) Zero-shot downstream performance comparison. All models are trained for 100K steps using FineWeb_edu data without data repetition.
Models Recipe Layers Perf. Medium rand init 24 47.74 rand init 16 46.84 Ours 16 48.08 Large rand init 32 49.44 rand init 16 48.17 Ours 16 49.75
Baselines.
We compare GPT-2 model variants (variants are models with lesser layers compared to their vanilla configuration) trained using $\mathsf{Inheritune}$ against two types of baselines: (1) larger reference models with more layers, and (2) similar-sized models with the same number of layers trained from scratch (randomly initialized). Additionally, we compare $\mathsf{Inheritune}$ against efficient training baselines that leverage pre-trained model initialization. These include stacking Gong et al. (2019); J. Reddi et al. (2023), hybrid stacking (which initializes the model using layers from the reference model), and half-width initialization (where all layers are retained, but the hidden dimension and number of attention heads are halved and initialized from the reference model). Finally we briefly compare $\mathsf{Inheritune}$ with knowledge distillation (Hinton et al., 2015). Detailed descriptions of these baselines are provided in the supplementary material Section B.
@@
Table 1: $\mathsf{Inheritune}$ outperforms baseline zero-shot initialization and efficient training techniques. Comparison of pre-training validation loss for GPT-2 xLarge, GPT-2 Large and GPT-2 Medium variants. $\mathsf{Inheritune}$-derived models consistently achieve lower loss compared to models initialized with stacking, hybrid stacking, and half-width techniques.

### v4

@@ Abstract
Large Language Models (LLMs) are known for their performance, but we uncover a significant structural inefficiency: a phenomenon we term attention collapse. In many pre-trained decoder-style LLMs, the attention matrices in deeper layers degenerate, collapsing to near rank-one structures. These underutilized layers, which we call lazy layers, are redundant and impair model efficiency. To address this, we introduce $\mathsf{Inheritune}$, a simple yet powerful training recipe designed to build smaller, stronger language models. Inheritune initializes a compact model by inheriting the potent early layers from a larger pre-trained model and then progressively trains and expands it. Our experiments on various models, including the GPT-2 family, demonstrate that models trained with $\mathsf{Inheritune}$ can match or even surpass the performance of their larger counterparts, despite having significantly fewer layers. This work presents a novel path toward model compression by design, enabling the creation of compact, yet highly performant language models. Code is available at https://github.com/sanyalsunny111/LLM-Inheritune.
@@ 1 Introduction
Large Language Models (LLMs) are composed of stacks of decoder-style transformer blocks (Vaswani et al., 2017). As the model grows in size, the model capacity and performance typically improve (Kaplan et al., 2020; Hoffmann et al., 2022). A substantial fraction of the total parameters is devoted towards adding more transformer blocks to increase the depth. Each block or layer in the stack refines the representations learned by the previous blocks, allowing the model to develop a nuanced understanding of the input data.
A transformer block primarily consists of a self-attention module and a feed-forward network (FFN, also referred to as an MLP). Among these, the causal self-attention mechanism (hereafter referred to as attention) is arguably the most critical component. It enables the model to combine token embeddings as a weighted linear sum of attention scores, effectively capturing long-range dependencies and contextual relationships within text. However, as models become deeper, they often exhibit a phenomenon known as attention degeneration, characterized by a collapse in the rank of the attention matrices. While prior studies have analyzed rank collapse in simplified transformer settings (Dong et al., 2021; Noci et al., 2022; He et al., 2023), this phenomenon has not, to our knowledge, been systematically explored in standard decoder-only LLMs. A formal discussion of attention degeneration is provided in Section 2. In this paper, we conduct a detailed empirical analysis of attention degeneration in the GPT-2 family of LLMs (Radford et al., 2019), including GPT-2 Medium (355M), GPT-2 Large (770M), and GPT-2 XLarge (1.5B). Our analysis reveals that many deeper layers in these models exhibit predominantly rank-1 attention matrices across most attention heads within a layer. This suggests that the attention mechanism loses its discriminative ability among tokens and instead performs near-uniform averaging across the sequence. We refer to layers in which all attention matrices degenerate to near rank-1 as lazy layers (a more formal definition is provided in Definition G.1). In the supplementary material, we further extend this analysis to billion-sized LLMs, including LLaMA-3 8B (refer Figure 11), Falcon-7B (refer Figure 21), OLMo-1B, Cerebras GPT 2.7B and LLaMA-3 3B (refer Figure 22), to highlight that attention collapse persists even in several modern architectures.
@@
Motivated by the new finding through our novel analysis we aim to develop performant small base language models (LMs) utilizing weights from inefficient larger base LMs without losing pre-train performance (measured by train/validation loss). A base LM is a decoder-style model trained solely for next-token prediction without additional enhancements like instruction tuning or reinforcement learning with human feedback (RLHF). Our proposal is straightforward, we start by initializing our smaller LM (target) using the first few blocks from a large pre-trained LM (reference). We then train the target model for a specified number of steps. After this initial training, we incrementally grow the target model by adding more blocks, continuing the training process until it matches or surpasses the pre-train validation loss (also val loss) of the reference model. During the growth phase, the newly added blocks can be initialized with lazy layers of the reference LM. We refer to this simple yet effective training approach as $\mathsf{Inheritune}$.
(a) GPT-2 Medium
(b) GPT-2 Large
(c) GPT-2 XLarge
@@
    Novel analysis of attention degeneration in standard decoder LLMs. We empirically investigate attention degeneration in standard decoder style LLMs. Our analysis shows that rank-collapse in attention matrices, revealing a significant structural inefficiency in the attention mechanism of standard LLMs in deeper layers (refer Figure 1). This degeneration gives rise to what we refer to as lazy layers.
@@
    Introduction of our training recipe Inheritune. Building on our analysis we observe that deep LMs often fail to fully utilize their effective depth. To address this inefficiency, we propose $\mathsf{Inheritune}$— a simple yet effective training recipe for developing smaller LMs without losing pre-training performance. This method involves inheriting a few early blocks from a much larger reference model and progressively growing and training the smaller model. We validate the effectiveness of $\mathsf{Inheritune}$ through extensive experiments on GPT-2 XLarge (1.5B), GPT-2 Large (770M), and GPT-2 Medium (355M) models, trained primarily on the OpenWebText dataset and additionally on FineWeb data.
@@
    Evaluation against multiple baselines. Models trained using $\mathsf{Inheritune}$ consistently outperform a wide range of baselines, including much larger models trained from scratch (see Figure 5 and Figure 7), as well as same sized models trained from scratch for twice as many steps (extended training; see Figure 6). We further compare against warm-started baselines (models initialized with pre-trained weights rather than random initialization  (Ash and Adams, 2020)) as shown in Table 2.
@@ 2 Attention Collapse and the Emergence of Lazy Layers in LLMs
A vanilla transformer-based model consists of $L$ transformer blocks (layers). The model operates on an input sequence $X\in\mathbb{R}^{T\times e}$, where $T$ denotes the sequence length (number of tokens), and $e$ represents the embedding dimension or model hidden size. The output of each layer $l$ is denoted as $X^{(l)}\in\mathbb{R}^{T\times e}$. Each transformer block primarily consists of two modules: a self-attention block and a feed-forward network (FFN). The self-attention mechanism enables the model to weight the relevance of different tokens in the sequence relative to each other. Specifically, for a single attention head, the attention computation is defined as equation 1.
$\text{Attention}(Q,K,V)=\underbrace{\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)}_{\textbf{Attention matrix: $A(X)$}}V$ (1)
@@
where $\tau\in(0,1)$ represents the proportion of variance that must be captured by the top $k$ singular values. A lower value of $k^{*}$ indicates stronger rank collapse. In this work, we set $\tau=0.90$.
In Figure 1, we present the layer-wise analysis of rank of GPT-2 models. For this analysis, we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText with 4M tokens, each with a sequence length of $T=100$ tokens across all attention heads within each layer. We then define the average approximate rank for each head and layer as $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$. Subsequently, we aggregate this metric per layer by taking the maximum rank across heads: $\mathrm{MaxRank}^{(l)}=\max_{h}\{\mathrm{Rank}^{(h,l)}\}$. As demonstrated in Figure 1, $\text{MaxRank}^{(l)}$ reveals that many deeper layers exhibit attention matrices that are predominantly near rank-1. We highlight that this rank collapse occurs in GPT-2 Medium, Large, and XLarge models, which are widely used modern LLMs thereby extending the limited findings of Dong et al. (2021) and Noci et al. (2022). We further visualize attention matrices from representative potent and lazy layers of GPT-2 XLarge in Figure 12. Overall, the degeneration of attention matrices in deeper layers provides quantitative evidence for the existence of lazy layers. Specifically, we observe that some deeper layers exhibit a near-complete rank collapse of attention matrices across all heads, suggesting potentially reduced representational capacity and less effective token mixing in these layers.
We provide an extended discussion of attention collapse in Appendix B. We analyze five modern LLMs for attention collapse (also refer Appendix G). We further assess the robustness of our findings by conducting collapse analyses on different datasets and by using an alternative metric, namely the mass of attention matrices. Finally, we perform an ablation over a range of values of $\tau$ (refer Appendix 4.3) to study the sensitivity of the rank-based analysis with respect to $\tau$.
@@ 2.1 The Functional Ineffectiveness of Lazy Layers
(a) Rank analysis of GPT-2 Small (12 layer).
(b) Performance of 4 layer GPT-2 small variants.
@@
Figure 2: Higher-rank (potent) layers transfer better. (Left, 2(a)) Layer-wise $\mathrm{MaxRank}^{(l)}$ of a pre-trained 12L GPT-2 Small. (Right, 2(b)) Validation loss of 4 layer variants initialized with potent layers (AvgRank $\approx 8.4-9.5$) vs. lazy layers (AvgRank $\approx 1.2$) and random weights, after training for 100K steps. Models initialized with lazy layers mirrors the model with random initialization. Training curves are smoothed for visual clarity.
@@
In the first set of experiments, we trained a vanilla GPT-2 small (125M) model with 12 layers for 100K steps on the OpenWebText dataset. We then performed the rank analysis described earlier, with results presented in Figure 2. Specifically, we aggregated the approximate ranks over groups of contiguous layers using $\text{AvgRank}=\frac{1}{L}\sum_{l=1}^{L}\text{MaxRank}^{(l)}$, where $L$ is the number of layers in each group. Subsequently, we trained three GPT-2 small variants [^1] A variant shares the same configurations as the reference model but has fewer layers. for 100K steps, each initialized with a different contiguous block of four layers from the trained vanilla GPT-2 small model: (a) layers 1–4, with $\text{AvgRank}=8.40$; (b) layers 5–8, with $\text{AvgRank}=9.48$; and (c) layers 9–12, with $\text{AvgRank}=1.22$. The last model is initialized with lazy layers. For comparison, we also trained another GPT-2 small variant with random initialization for 100K steps. All models were trained on the OpenWebText dataset. As shown in Figure 2, the model initialized with layers from the vanilla GPT-2 small model having higher AvgRank demonstrated the best performance (i.e., lowest final validation loss). Additionally, we observed that the model initialized with lazy layers performed very similarly to the model with random initialization suggesting that lazy layers contain minimal transferable knowledge. The results are also summarized in Table 4.
(a) GPT-2 Medium (12 layer) variants.
(b) GPT-2 Large (18-layer) variants.
@@
Figure 3: Initializing 12-layer and 18-layer variants of GPT-2 Medium and GPT-2 Large with deeper lazy layers exhibiting degenerated attention results in performance comparable to random initialization. In contrast, initializing with early (high-rank) potent layers leads to substantially better convergence and generalization. Training curves are smoothed for visual clarity.
@@
For the second set of experiments we utilized larger models namely GPT-2 Medium and GPT-2 Large both similarly trained for 100K steps using OpenWebText. Here we initialized a 12-layer GPT-2 Medium variant and an 18-layer variant of GPT-2 Large using lazy layers extracted from pre-trained 24-layer GPT-2 Medium and 36-layer GPT-2 Large models. We then trained these GPT-2 variants on the same dataset for 10K steps. For comparison, we conducted two baseline experiments where the GPT-2 variants were initialized either with the first half of transformer layers (potent layers with high AvgRank) and with random initialization. As shown in Figure 3, models initialized with lazy layers demonstrate poor transferability, performing similarly to models with random initialization. This provides additional evidence that lazy layers with fully degenerated attention, fails to learn meaningful representations.
Theoretical Analysis.
Finally, we analyze the implications of attention rank collapse on model training from a theoretical perspective (refer Section H). Our key insight is that rank-collapsed attention head(s) impede learning by inducing vanishing gradients, effectively suppressing updates to the associated $W_{Q}$ and $W_{K}$.
@@ 3 Inheritune: Our Proposed Training Recipe
6:while $\mathcal{M}_{\text{tgt}}$ performance << $\mathcal{M}_{\text{ref}}$ performance on $\mathcal{D}_{\text{val}}$ do
@@
Figure 4: Overview of the Inheritune training recipe using a 24-Layer GPT-2 Medium model example. A smaller target model is initialized using early layers from a larger, pre-trained reference model. The target model goes multiple rounds of training while inheriting contiguous layers until it matches/outperforms the reference model. The intensity of the red color in layers correlates with $\mathrm{MaxRank}^{(l)}$.
@@
  1. 1.
  2. 2.
  3. 3.
    Grow: If needed, increase $\mathcal{M}_{\text{tgt}}$’s size by adding the next few contiguous layers and repeat steps 1-2 until desired performance is achieved.
@@ 4 Experiments
We evaluate $\mathsf{Inheritune}$ through a comprehensive set of experiments using several GPT-2 models: a 48-layer GPT-2 XLarge (1.5B), a 36-layer GPT-2 Large (770M), a 32-layer GPT-2 Large (668M), and a 24-layer GPT-2 Medium (355M) (Radford et al., 2019). Table 6 provides detailed specifications of all model configurations used in our experiments.
We use two training datasets: OpenWebText Gokaslan and Cohen (2019) with 10B tokens and FineWeb (education subset) (Penedo et al., 2024) also with 10B tokens. Our experimental setup closely follows prior work (Liu et al., 2023; Sanyal et al., 2024). For models trained on OpenWebText, we report validation loss (log perplexity), while for models trained on FineWeb, we report training loss (also log perplexity).
Additionally, for models trained on FineWeb, we conduct zero-shot downstream evaluations using the lm-evaluation-harness (Gao et al., 2024) across five standard benchmarks: ARC-Easy (ARCE; Clark et al., 2018), LAMBADA (Paperno et al., 2016), SciQ (Welbl et al., 2017), HellaSwag (Zellers et al., 2019), and PIQA (Bisk et al., 2020). Finally, we perform a detailed ablation study by initializing each submodule within a transformer block in isolation and training it for 100K steps to identify which component contributes most to performance. All training curves are smoothed for visual clarity.
(a) GPT-2 Medium
(b) GPT-2 Large
(c) GPT-2 XLarge
Figure 5: Models derived using Inheritune converge faster and match the final validation loss of the full-sized model, despite having much fewer layers. Comparison of $\mathsf{Inheritune}$ trained models (24-layer GPT-2 XLarge variant, 18-layer GPT-2 Large variant, 16-layer GPT-2 Medium variant) against their full-sized counterparts and same sized variants trained from scratch. All models are trained for 100K steps using OpenWebText data.
(a) GPT-2 Medium
(b) GPT-2 Large
(c) GPT-2 XLarge
Figure 6: Models trained with Inheritune match the validation loss of same sized models trained from scratch for twice as many steps. We compare $\mathsf{Inheritune}$-trained models (24-layer GPT-2 XLarge, 18-layer GPT-2 Large, and 16-layer GPT-2 Medium) against their same-sized counterparts trained from scratch for twice the number of steps. $\mathsf{Inheritune}$ models are trained for 100K steps, while the baseline models (trained from scratch) are trained for 200K steps, on OpenWebText data.
We provide experimental details of our proposed training recipe $\mathsf{Inheritune}$ using a GPT-2 Medium model as an example; similar procedure was applied to train other models. A visualization of the training recipe is presented in Figure 4. Our recipe for applying $\mathsf{Inheritune}$ involves the following steps.
@@
    Reference Model: We train a vanilla 24-layer GPT-2 Medium model (reference model) on $\mathcal{D}_{\text{train}}$ for 100K steps and evaluate its validation loss ( log perplexity) on $\mathcal{D}_{\text{val}}$. This establishes our benchmark validation loss.
@@
    Iterative Refinement: If the smaller model’s performance is inferior to the reference model, then we incrementally increase its size by adding additional layers and repeat steps 2-3 until we achieve parity with the reference model’s validation loss.
@@
(a) GPT-2 Large
(b) GPT-2 Medium
Figure 7: Models derived using Inheritune without data repetition converge faster and match the final validation loss of the full-sized model despite using lesser layers. Additionally, the model trained using $\mathsf{Inheritune}$ demonstrates data efficiency, achieving a lower validation loss in fewer steps compared to its full-sized and half-sized counterparts.
Models Recipe Layers ARCE PIQA SciQ Hellaswag Lambada Average
GPT-2 Medium rand init 24 51.05 61.81 74.8 30.79 20.28 47.74
rand init 16 49.92 61.92 73.3 29.56 19.54 46.84
Ours 16 51.26 61.81 73.8 30.55 23 48.08
GPT-2 Large rand init 32 52.48 64.58 75.3 32.65 22.2 49.44
rand init 16 50.34 63.11 75 30.86 21.56 48.17
Ours 16 52.9 63.55 76.1 32.14 24.06 49.75
Table 1: Models trained with Inheritune achieve comparable average zero-shot downstream performance to their larger reference models and surpass same-sized counterparts trained from scratch. Downstream evaluations are performed on models pre-trained with the FineWeb dataset (see Figure 7). Performance is measured using accuracy (acc) and normalized accuracy (acc-norm) metrics, following the Open LLM Leaderboard protocol Beeching et al. (2023). We have highlighted the best average scores in bold.
Baseline-I.
We compare GPT-2 model variants (i.e., models with fewer layers than their vanilla configurations) trained using $\mathsf{Inheritune}$ against the following baselines:
  1. 1.
    Larger reference models with more layers, trained from scratch (random initialization).
  2. 2.
    Same-sized models with the same number of layers, trained from scratch.
  3. 3.
    Extended training baselines, same-sized models trained from scratch for twice as many steps compared to models trained with $\mathsf{Inheritune}$.
Baseline-II.
Additionally, we compare $\mathsf{Inheritune}$ against warm-started baselines (Ash and Adams, 2020). Warm starting in neural network training refers to initializing a model’s parameters with weights from a previously trained model, rather than starting from random initialization (a “cold start”). This setup enables a fair comparison with our approach, which also leverages prior learned representations. The warm-started baselines include stacking (Gong et al., 2019; J. Reddi et al., 2023), hybrid stacking (which initializes model layers directly from a larger reference model), and half-width (which retains all layers but reduces both the hidden dimension and number of attention heads by half, using weights from the reference model). Finally, we briefly compare $\mathsf{Inheritune}$ with knowledge distillation (Hinton et al., 2015); results are provided in the supplementary material (see Figure 17). Detailed descriptions of all baselines are presented in Section D of the supplementary material.
@@
Table 2: Inheritune outperforms warm-started baselines (Baseline-II). Comparison of pre-training validation loss for GPT-2 XLarge, GPT-2 Large, and GPT-2 Medium variants trained with Stacking, Hybrid Stacking, Half-Width, and $\mathsf{Inheritune}$ recipes. All baselines are warm-started, i.e., initialized with pre-trained weights rather than random initialization. Across model scales, $\mathsf{Inheritune}$ consistently achieves lower validation loss than the majority of warm-started baselines. The lowest validation loss (lower is better) is highlighted in bold.